# Detecting Collective Excitations in Self-Gravitating Bose-Einstein Condensates via Faraday Waves

### v1

@@ Abstract
We propose Faraday waves as a probe for collective excitations in self-gravitating Bose-Einstein condensates (SGBECs), driven by periodic modulation of the $s$-wave scattering length. Linear stability analysis of the driven Gross-Pitaevskii-Newton equations reveals that parametric instability follows a Mathieu-like equation, with Faraday waves emerging resonantly when half the driving frequency matches the collective excitation frequency of the SGBEC. This framework yields a stability phase diagram demonstrating the coexistence and interplay between the intrinsic Jeans instability of self-gravitating systems and the unstable tongues of parametric resonance. The phase diagram shows that reducing the driving frequency fragments and narrows parametric resonance structures while expanding the Jeans-dominated regime. We further identify a thermal-gravity competition in the Faraday wavevector, manifested as a sharp transition when thermal energy balances mean-field interaction at low driving frequencies. We numerically simulate Faraday wave formation and dynamics within the SGBEC, including the effects of dissipation; simulations reveal that Faraday wave patterns exhibit high sensitivity to Jeans frequency, transitioning from parametric resonance to Jeans collapse as gravitational strength increases.
@@ I Introduction
Gravity, historically the first fundamental interaction to be characterized, presents unresolved challenges in quantum regimes. Persistent uncertainty surrounds its fundamental nature: whether gravity is quantum or classical Oppenheim (2023); Kryhin and Sudhir (2025). This uncertainty persists because the extreme weakness of gravitational interactions prevents definitive experimental tests of quantum gravitational behavior. Consequently, analog quantum systems that simulate gravitational phenomena under controlled conditions offer a vital pathway forward. These systems engineer long-range attractive interactions that simulate key aspects of gravity. This approach provides crucial insights into collective quantum behavior under gravitational-like potentials. It also helps distinguish intrinsic gravitational features from those of other fundamental interactions Barcelo et al. (2011); Liang et al. (2024a); Marletto and Vedral (2025).
Within this context, self-gravitating quantum many-body systems present a particularly intriguing but challenging subject Chavanis (2002); Andersson and Schmidt (2009); Großardt (2016); Ourabah (2020, 2023); Asakawa et al. (2024). Yet even foundational aspects, such as their collective excitation spectra, remain topics of debate Chavanis (2017); Akbari-Moghanjoughi (2024). Recent advances in ultracold atomic physics, however, offer an excellent solution. The exceptional purity and unprecedented tunability of ultracold atomic gases make them an ideal platform for quantum simulation Morsch and Oberthaler (2006); Bloch et al. (2008). Crucially, the collective behavior of self-gravitating systems, akin to ensembles with long-range attractive interactions, can be engineered in these gases. This can be experimentally achieved by inducing effective gravitational attraction through the application of non-resonant, intense optical dipole forces from laser beams O’dell et al. (2000); Giovanazzi et al. (2001); Choi (2002), or through two contra-propagating quasiresonant lasers Chalony et al. (2013).
@@
Building on these experimental capabilities, we focus specifically on self-gravitating Bose-Einstein condensates (SGBECs). SGBECs are governed by the Gross-Pitaevskii-Newton (GPN) equations, which couple the nonlinear Schrödinger equation describing the condensate wavefunction to the Poisson equation for the self-consistent gravitational potential Paredes et al. (2020). The GPN framework has enabled extensive analysis of SGBECs, including nonlinear dynamics Cartarius et al. (2008), equilibrium configurations Guzman and Urena-Lopez (2006), and quantized vortex formation Asakawa and Tsubota (2024). Beyond their role as laboratory analogs, SGBECs are of significant interest in astrophysics and cosmology Hu et al. (2000); Boehmer and Harko (2007); Jones and Bernstein (2001); Schroven et al. (2015); Chavanis (2020); Guo et al. (2021); Kirejczyk et al. (2022); Sivakumar et al. (2025). They serve as models for ultralight bosonic dark matter candidates Chavanis (2011a, b, 2014); Suárez et al. (2014); Chavanis (2021); Nikolaieva et al. (2021, 2023), potentially explaining core structures in galactic halos, and for boson stars Kolb and Tkachev (1993); Chavanis and Harko (2012); Dmitriev et al. (2021); Levkov et al. (2017); Dmitriev et al. (2024), hypothetical compact objects composed of scalar fields. Understanding the dynamics and collective excitations of SGBECs is thus crucial for probing fundamental physics across scales.
@@
To probe the collective excitations of SGBECs, the generation and observation of Faraday waves provide a powerful diagnostic tool Hernández-Rajkov et al. (2021); Cominotti et al. (2022). Faraday waves are striking standing wave patterns that emerge on a fluid surface subjected to parametric driving, i.e., when an external control parameter (like vertical oscillation amplitude) is varied periodically. First documented by Michael Faraday in 1831 Faraday (1831), this phenomenon represents a spontaneous breaking of both temporal and spatial symmetry. Significantly, Faraday waves have been investigated in diverse quantum systems Staliunas et al. (2002, 2004); Nicolin et al. (2007); Bhattacherjee (2008); Nath and Santos (2010); Nicolin and Raportaru (2010); Pollack et al. (2010); Nicolin (2011); Capuzzi et al. (2011); Balaž and Nicolin (2012); Łakomy et al. (2012); Raportaru (2012); Balaž et al. (2014); Abdullaev et al. (2015); Sudharsan et al. (2016); Nguyen et al. (2019); Vudragović and Balaž (2019); Abdullaev et al. (2019, 2019); Turmanov et al. (2020); Liu and Tu (2020); Cheng and Shi (2021); Zhang et al. (2020); Maity et al. (2020); Kwon et al. (2021); Zhang et al. (2022); Shukuno et al. (2023); Brito et al. (2023); Liang et al. (2024b); Wan et al. (2024); del Río-Lima et al. (2024); Wang et al. (2025), including their first experimental demonstration in atomic BECs by Engels et al. Engels et al. (2007) and subsequent studies in dipolar condensates Łakomy et al. (2012); Turmanov et al. (2020), two-component BECs Chen et al. (2019); Wang et al. (2025), spin-orbit-coupled BECs Zhang et al. (2022) and Rydberg-dressed BECs Shi et al. (2023). These theoretical analyses and experiments establish that Faraday waves directly probe the underlying collective excitation spectrum of the quantum fluid. Therefore, driving Faraday waves in SGBECs offers a novel and controlled framework for investigating the collective dynamics, particularly the excitation spectrum of quantum fluids dominated by long-range attractive interactions.
@@ II.1 The Gross-Pitaevskii-Newton equation
The dynamics of self-gravitating Bose-Einstein condensates (SGBECs) is governed by the coupled Gross-Pitaevskii-Newton(Possion) (GPN or GPP) equations:
${\rm i}\hbar\frac{\partial\Psi}{\partial t}$ $=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi,$ (1)
$\nabla^{2}\Phi$ $=4\pi Gm|\Psi|^{2},$
@@
[…]
Without gravity ($\Phi=0$), Eq. (1) reduces to the standard Gross-Pitaevskii equation; neglecting the mean-field term ($g=0$) yields the Schrödinger-Newton equation.
@@
where $\gamma$ is the damping coefficient. The term $-\beta^{-1}\ln|\Psi|^{2}\Psi$ accounts for finite-temperature effects, where $\beta^{-1}=-k_{B}T$ ($k_{B}$ is Boltzmann’s constant, $T$ is temperature), and the last term describes phenomenological dissipation. Eq. (1) is recovered when $\beta^{-1}=0$ and $\gamma=0$.
II.2 Linear stability analysis
Faraday waves in BECs are excited by periodic modulation of either the external potential or the coupling strength $g(t)$ Engels et al. (2007). Here we modulate $g(t)$ as
$g(t)=g_{0}\left(1+\alpha\cos\omega t\right),$ (4)
@@ II.2 Linear stability analysis
where $g_{0}=4\pi\hbar^{2}a_{s}/m$ is the unmodulated coupling strength, $\omega$ is the driving frequency, and $\alpha$ is the modulation depth. We consider a cigar-shaped condensate tightly confined in the transverse ($xy$) plane, with driving along the longitudinal ($z$) direction. When the Faraday wavelength $k^{-1}$ is much smaller than the condensate radius $R_{\rm TF}$, the external potential $V_{\rm ext}$ is negligible.
To analyze stability, we decompose $\Psi=\Psi_{B}\Psi_{P}$, where $\Psi_{B}(\mathbf{r}_{\perp})$ is the transverse ground state and $\Psi_{P}(z,t)$ is the longitudinal perturbation. The equilibrium state satisfies Hernández-Rajkov et al. (2021); Hai et al. (2008)
$g_{0}|\Psi_{B}|^{2}+m\Phi_{B}-\beta^{-1}\ln|\Psi_{B}|^{2}=E_{B},$ (5)
with $\nabla^{2}\Phi_{B}=4\pi mG|\Psi_{B}|^{2}$, where $E_{B}$ is the equilibrium energy eigenvalue. The unperturbed dynamics follows
${\rm i}\hbar\frac{\partial\Psi_{B}}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}_{\perp}+E_{B}\right]\Psi_{B}.$ (6)
Substituting $\Psi=\Psi_{B}\Psi_{P}$ into Eq. (3) and using Eq. (6) yields the perturbation equation:
${\rm i}\hbar\frac{\partial\Psi_{P}}{\partial t}=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}\Psi_{P}}{\partial z^{2}}+g(t)|\Psi_{B}|^{2}\left(|\Psi_{P}|^{2}-1\right)\Psi_{P}+m(\Phi-\Phi_{B})\Psi_{P}-\beta^{-1}\ln|\Psi_{P}|\Psi_{P}-{\rm i}\hbar\gamma\Psi_{P}.$ (7)
Linearizing around $\Psi_{P}=1$ for small perturbations $\Psi_{P}(z,t)=1+W(t)\cos(kz)$ with $W(t)=U(t)+{\rm i}V(t)$ gives coupled equations:
$\hbar\dot{V}$ $=-\epsilon_{k}U-2g(t)|\Psi_{B}|^{2}U+\frac{2m\Omega_{J}^{2}}{k^{2}}U+\beta^{-1}U+\hbar\gamma V,$ (8)
$\hbar\dot{U}$ $=\epsilon_{k}V+\hbar\gamma U,$
where $\epsilon_{k}=\hbar^{2}k^{2}/(2m)$ is the free-particle energy, and $\Omega_{J}=\sqrt{4\pi Gm|\Psi_{B}|^{2}}$ is the Jeans frequency. Combining these yields a damped Mathieu equation:
$\frac{{\rm d}^{2}U}{{\rm d}t^{2}}+2\gamma\frac{{\rm d}U}{{\rm d}t}+\left[\Omega_{k}^{2}+\frac{\alpha g_{0}|\Psi_{B}|^{2}}{m}k^{2}\cos\omega t\right]U=0,$ (9)
@@
$\Omega_{\rm Bog}^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+c_{s}^{2}k^{2},\quad c_{s}^{2}=\frac{g_{0}n}{m}-\frac{\beta^{-1}}{m},$ (11)
where $n=|\Psi_{B}|^{2}$ is the equilibrium density, and $c_{s}$ is the sound speed. The term $g_{0}n/m$ arises from mean-field interactions, and $-\beta^{-1}/m$ from finite-temperature effects. In the zero-temperature limit ($\beta^{-1}=0$), $c_{s}^{2}=g_{0}n/m$.
Fig. 1 shows the collective excitation spectrum of SGBECs. For large $k$ ($k\xi\gg 1$), where $\xi=\hbar/\sqrt{2mg_{0}n}$ is the healing length, $\Omega_{k}$ approaches the Bogoliubov dispersion. Near $k_{J}\xi\equiv\sqrt{\sqrt{1+\Omega^{2}_{J}/\Omega^{2}_{0}}-1}$ (with $\Omega_{0}=g_{0}n/\hbar$), the Jeans instability causes $\Omega_{k}^{2}<0$, indicating exponential growth. Here $k_{J}$ is the critical wavenumber defining the quantum Jeans scale Chavanis (2020).
Figure 1: (Color online) Collective excitation spectrum of SGBECs. Parameters: $\Omega_{0}=g_{0}n/\hbar$, $\xi=\hbar/\sqrt{2mg_{0}n}$. The region $k<k_{J}$ indicates Jeans instability.
Using the dimensionless time $\tau=\omega t/2$, Eq. (9) becomes
$\frac{d^{2}U}{d\tau^{2}}+\bar{\gamma}\frac{dU}{d\tau}+\left[\lambda-2q\cos(2\tau)\right]U=0,$ (12)
where $\bar{\gamma}=4\gamma/\omega$, and
$\lambda=\frac{4\Omega_{k}^{2}}{\omega^{2}},\quad q=-\frac{2\alpha g_{0}n}{m}\frac{k^{2}}{\omega^{2}}.$ (13)
The parameter $\lambda$ encodes the excitation spectrum of the self-gravitating quantum fluid. For $\gamma=0$, Eq. (12) reduces to the canonical Mathieu equation, with well-established stability properties McLachlan (1947); Abramowitz and Stegun (1965); Wang and Guo (1989). We analyze its phase diagram and derive the Faraday wavevector in the following section.

### v2

@@ Abstract
We propose Faraday waves as a probe for collective excitations in self-gravitating Bose-Einstein condensates (SGBECs), driven by periodic modulation of the $s$-wave scattering length. Linear stability analysis of the driven Gross-Pitaevskii-Newton equations reveals that parametric instability follows a Mathieu-like equation, with Faraday waves emerging resonantly when half the driving frequency matches the collective excitation frequency of the SGBEC. This framework yields a stability phase diagram that maps the competitive interplay between the unstable tongues of parametric resonance and the intrinsic Jeans instability. The diagram reveals that increasing the driving frequency compresses the Jeans-unstable region and allows well-separated parametric resonance tongues to dominate, thereby creating a clear regime for observing Faraday waves. Conversely, lowering the driving frequency expands the domain of Jeans instability, which fragments and overwhelms the parametric resonance structures. We numerically simulate Faraday wave formation and dynamics within the SGBEC, including the effects of dissipation; simulations reveal a characteristic transition from parametric-resonance-driven Faraday waves to gravity-dominated Jeans collapse as the Jeans frequency increases.
@@ I Introduction
Gravity, the first fundamental interaction to be characterized, remains the least understood in quantum regimes Oppenheim (2023); Kryhin and Sudhir (2025). The extreme weakness of gravitational interactions prevents direct experimental probes of quantum gravitational phenomena in tabletop experiments. Analog quantum systems that emulate gravitational phenomena through engineered long-range attractive interactions provide a valuable alternative for studying quantum many-body behavior under gravitational-like potentials. These systems offer insights into collective quantum dynamics that may parallel those in real gravitational systems, serving as testbeds for theoretical ideas about quantum behavior in gravitational settings Barcelo et al. (2011); Liang et al. (2024a); Marletto and Vedral (2025).
Within this context, self-gravitating quantum many-body systems have attracted growing interest, particularly as analog platforms for exploring gravitational phenomena in controllable quantum settings Chavanis (2002); Andersson and Schmidt (2009); Großardt (2016); Ourabah (2020, 2023); Asakawa et al. (2024). These systems are of fundamental importance for testing quantum-gravitational phenomena and have wide-ranging implications for modeling astrophysical scenarios. Notably, even foundational issues such as the form of the collective excitation spectrum remain contentious. For instance, a hydrodynamic approach yields a classical-like spectrum Chavanis (2017), while a direct perturbative treatment of the quantum equations predicts ultraviolet divergence Akbari-Moghanjoughi (2024). Resolving these theoretical discrepancies requires experimental access to the excitation spectrum of such systems. Ultracold atomic gases, with their exceptional purity and tunability, offer an ideal platform for quantum simulation of these phenomena Morsch and Oberthaler (2006); Bloch et al. (2008). Through techniques such as optical dipole forces or laser-induced interactions, which can generate long-range attractive interactions analogous to gravity O’dell et al. (2000); Giovanazzi et al. (2001); Choi (2002); Chalony et al. (2013), they enable the emulation of self-gravitating quantum fluids and the study of collective behavior under controlled, gravitational-like potentials.
@@
Building on these experimental capabilities, we focus specifically on self-gravitating Bose-Einstein condensates (SGBECs). SGBECs are governed by the Gross-Pitaevskii-Newton (GPN) equations, which couple the nonlinear Schrödinger equation describing the condensate wavefunction to the Poisson equation for the self-consistent gravitational potential Paredes et al. (2020). The GPN framework has enabled extensive analysis of SGBECs, including nonlinear dynamics Cartarius et al. (2008), equilibrium configurations Guzman and Urena-Lopez (2006), and quantized vortex formation Asakawa and Tsubota (2024). Beyond their role as laboratory analogs, SGBECs are of significant interest in astrophysics and cosmology Hu et al. (2000); Boehmer and Harko (2007); Jones and Bernstein (2001); Schroven et al. (2015); Chavanis (2020); Guo et al. (2021); Kirejczyk et al. (2022); Shukla et al. (2024); Sivakumar et al. (2025). They serve as models for ultralight bosonic dark matter candidates Chavanis (2011a, b, 2014); Suárez et al. (2014); Verma et al. (2021); Chavanis (2021); Nikolaieva et al. (2021, 2023), potentially explaining core structures in galactic halos, and for boson stars Kolb and Tkachev (1993); Chavanis and Harko (2012); Dmitriev et al. (2021); Verma et al. (2022); Levkov et al. (2017); Dmitriev et al. (2024), hypothetical compact objects composed of scalar fields. Understanding the dynamics and collective excitations of SGBECs is thus crucial for probing fundamental physics across scales.
@@
To probe the collective excitations of SGBECs, the generation and observation of Faraday waves provide a powerful diagnostic tool Hernández-Rajkov et al. (2021); Cominotti et al. (2022). Faraday waves are striking standing wave patterns that emerge on a fluid surface subjected to parametric driving, i.e., when an external control parameter (like vertical oscillation amplitude) is varied periodically. First documented by Michael Faraday in 1831 Faraday (1831), this phenomenon represents a spontaneous breaking of both temporal and spatial symmetry. Significantly, Faraday waves have been investigated in diverse quantum systems Staliunas et al. (2002, 2004); Nicolin et al. (2007); Bhattacherjee (2008); Nath and Santos (2010); Nicolin and Raportaru (2010); Pollack et al. (2010); Nicolin (2011); Capuzzi et al. (2011); Balaž and Nicolin (2012); Łakomy et al. (2012); Raportaru (2012); Balaž et al. (2014); Abdullaev et al. (2015); Sudharsan et al. (2016); Nguyen et al. (2019); Vudragović and Balaž (2019); Abdullaev et al. (2019, 2019); Turmanov et al. (2020); Liu and Tu (2020); Cheng and Shi (2021); Zhang et al. (2020); Maity et al. (2020); Kwon et al. (2021); Zhang et al. (2022); Shukuno et al. (2023); Brito et al. (2023); Liang et al. (2024b); Wan et al. (2024); del R´ıo-Lima et al. (2024); Wang et al. (2025), including their first experimental demonstration in atomic BECs by Engels et al. Engels et al. (2007) and subsequent studies in dipolar condensates Łakomy et al. (2012); Turmanov et al. (2020), two-component BECs Chen et al. (2019); Wang et al. (2025), spin-orbit-coupled BECs Zhang et al. (2022) and Rydberg-dressed BECs Shi et al. (2023). These theoretical analyses and experiments establish that Faraday waves directly probe the underlying collective excitation spectrum of the quantum fluid. Therefore, driving Faraday waves in SGBECs offers a novel and controlled framework for investigating the collective dynamics, particularly the excitation spectrum of quantum fluids dominated by long-range attractive interactions.
@@ II.1 The Gross-Pitaevskii-Newton equation
The dynamics of self-gravitating Bose-Einstein condensates (SGBECs) is governed by the coupled Gross-Pitaevskii-Newton (Poisson) (GPN or GPP) equations:
$\mathrm{i}\hbar\frac{\partial\Psi}{\partial t}$ $=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\mathrm{ext}}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi,$ (1a)
$\nabla^{2}\Phi$ $=4\pi Gm|\Psi|^{2}.$ (1b)
@@
${\rm i}\hbar\frac{\partial\Psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi-m^{2}G\int d\mathbf{r}^{\prime}\frac{|\Psi(\mathbf{r}^{\prime},t)|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}\Psi.$ (2)
The apparent singularity at $\mathbf{r}\to\mathbf{r}^{\prime}$ in the integral formulation is regularized by the finite extent of the wavefunction $|\Psi(\mathbf{r}^{\prime},t)|^{2}$, which acts as a natural cutoff. In practice, for a smooth, square-integrable wavefunction, this integral remains finite. The equivalence between the differential and integral forms can be rigorously established using distribution theory, where the Poisson equation $\nabla^{2}\Phi=4\pi Gm|\Psi|^{2}$ is interpreted in the weak sense, and the solution is given by the convolution with the Newtonian potential ${1}/{|\mathbf{r}-\mathbf{r}^{\prime}|}$, which is a fundamental solution of the Laplacian. Without gravity ($\Phi=0$), Eq. (1) reduces to the standard Gross-Pitaevskii equation; neglecting the mean-field term ($g=0$) yields the Schrödinger-Newton equation.
@@
where $\gamma$ is the damping coefficient. The term $\beta^{-1}\ln|\Psi|\Psi$ accounts for finite-temperature effects, where $\beta^{-1}=-k_{B}T$ ($k_{B}$ is Boltzmann’s constant, $T$ is temperature), and the last term describes phenomenological dissipation. Eq. (1) is recovered when $\beta^{-1}=0$ and $\gamma=0$.
II.2 Reduction to Quasi-One-Dimensional System
We consider a cigar-shaped condensate confined by a fully anisotropic harmonic potential $V_{\rm ext}(\mathbf{r})=\frac{1}{2}m[\omega_{\perp}^{2}(x^{2}+y^{2})+\omega_{z}^{2}z^{2}]$, where $\omega_{\perp}$ and $\omega_{z}$ represent the transverse and longitudinal confinement frequencies, respectively. To ensure the system maintains a quasi-one-dimensional character, we require strong transverse confinement relative to the longitudinal direction, i.e., $\omega_{\perp}\gg\omega_{z}$. Under this condition, the transverse degrees of freedom are energetically frozen into the ground state, while the longitudinal motion remains dynamic. This allows for the factorization of the wave function as:
$\Psi(\mathbf{r},t)=\psi_{B}(\mathbf{r}_{\perp})\psi_{P}(z,t),$ (4)
where $\psi_{B}(\mathbf{r}_{\perp})$ is the transverse ground state, and $\psi_{P}(z,t)$ describes the longitudinal dynamics of the condensate. The transverse ground state is well approximated by a Gaussian:
$\psi_{B}(\mathbf{r}_{\perp})=\frac{1}{\sqrt{\pi}a_{\perp}}e^{-(x^{2}+y^{2})/(2a_{\perp}^{2})},\quad a_{\perp}=\sqrt{\frac{\hbar}{m\omega_{\perp}}}.$ (5)
Here, $a_{\perp}$ is the transverse harmonic oscillator length, which characterizes the size of the condensate in the transverse direction.
Substituting Eq.(4) into Eq.(3) and integrating over transverse coordinates, we obtain the effective quasi-1D equation(see Appendix A for details):
${\rm i}\hbar\frac{\partial\psi_{P}}{\partial t}=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}\psi_{P}}{\partial z^{2}}+g_{\rm 1D}(t)|\psi_{P}|^{2}\psi_{P}+m\Phi_{\rm 1D}\psi_{P}-\beta^{-1}\ln|\psi_{P}|\psi_{P}-{\rm i}\hbar\gamma\psi_{P},$ (6)
where the effective 1D coupling is $g_{\rm 1D}(t)=g(t)/(2\pi a_{\perp}^{2})$ and $\Phi_{\rm 1D}$ is the quasi-1D gravitational potential. For small perturbations around the homogeneous background $\psi_{P}^{(0)}=\sqrt{n_{z}}$, where $n_{z}$ is the linear density (number of atoms per unit length), the self-gravity term can be approximated in the long-wavelength limit ($ka_{\perp}\ll 1$):
$m\Phi_{\rm 1D}\psi_{P}=-\frac{2Gm^{2}}{a_{\perp}}|\psi_{P}|^{2}\psi_{P}.$ (7)
This approximation is valid when the characteristic wavelength of density perturbations is much larger than the transverse confinement length scale $a_{\perp}$.
II.3 Linear stability analysis
Faraday waves in Bose-Einstein condensates can be excited through two primary methods: periodic modulation of the external trapping potential or time-dependent variation of the interaction strength via Feshbach resonance Engels et al. (2007). This work employs the latter approach, modulating the scattering length to produce a time-dependent quasi-1D coupling strength:
$g_{\rm 1D}(t)=g_{\rm 1D,0}\left(1+\alpha\cos\omega t\right),$ (8)
@@ II.3 Linear stability analysis
where $g_{\rm 1D,0}=g_{0}/(2\pi a_{\perp}^{2})=2\hbar^{2}a_{s}/(ma_{\perp}^{2})$ is the unmodulated quasi-1D coupling strength, $\omega$ is the driving frequency, and $\alpha$ is the modulation depth. We consider a cigar-shaped condensate tightly confined transversely, with driving applied along the longitudinal $z$-direction.We consider a cigar-shaped condensate tightly confined transversely, with driving applied along the longitudinal $z$-direction.
To analyze the stability against density modulations, we consider a plane wave perturbation with a general wavenumber $k$. The characteristic wavelength of such a modulation is $\lambda=2\pi/k$. The periodic driving will resonantly amplify a specific mode whose wavenumber $k_{F}$ (and corresponding wavelength $\lambda_{F}=2\pi/k_{F}$) satisfies the parametric resonance condition. When this Faraday wavelength $\lambda_{F}$ satisfies $\lambda_{F}\ll R_{\rm TF}$ (where $R_{\rm TF}$ is the Thomas-Fermi radius in the longitudinal direction), the external potential $V_{\rm ext}$ can be neglected, and the system can be treated as effectively homogeneous for the purpose of analyzing the Faraday wave formation.
The unperturbed state is homogeneous: $\psi_{P}^{(0)}=\sqrt{n_{z}}$, We introduce a perturbation:
$\psi_{P}(z,t)=\sqrt{n_{z}}\left[1+W(t)\cos(kz)\right],\quad W(t)=U(t)+{\rm i}V(t).$ (9)
[…]
$\hbar\dot{V}$ $=-\epsilon_{k}U-2g_{\mathrm{1D}}(t)\,n_{z}U+\frac{2m\Omega_{J,\mathrm{1D}}^{2}}{k^{2}}U+\beta^{-1}U-\hbar\gamma V,$ (10a)
$\hbar\dot{U}$ $=\epsilon_{k}V-\beta^{-1}V+\hbar\gamma U,$ (10b)
where $\epsilon_{k}=\hbar^{2}k^{2}/(2m)$ is the free-particle energy, and $\Omega_{J,{\rm 1D}}=\sqrt{4\pi Gm^{2}n_{2D}}$ is the quasi-1D Jeans frequency, with $n_{2D}=n_{z}/a_{\perp}$ being the effective areal density (number of atoms per unit area). The Jeans frequency derives its name from Sir James Jeans, who in 1902 first analyzed the gravitational instability of an infinite homogeneous medium in the context of astrophysics Jeans (1902). The Jeans criterion determines when self-gravity overcomes pressure support, leading to collapse Khlopov et al. (1985). In our quantum analog, $\Omega_{J,{\rm 1D}}$ represents the characteristic frequency at which self-gravity operates in the quasi-1D geometry, playing a role analogous to the classical Jeans frequency in astrophysical systems. For notational simplicity, we will hereafter denote $\Omega_{J,\mathrm{1D}}$ simply as $\Omega_{J}$ in the remainder of this paper.
Combining these equations gives a damped Mathieu equation:
$\frac{{\rm d}^{2}U}{{\rm d}t^{2}}+2\gamma\frac{{\rm d}U}{{\rm d}t}+\left[\Omega_{k}^{2}+\frac{\alpha g_{\rm 1D,0}n_{z}}{m}k^{2}\cos\omega t\right]U=0,$ (11)
@@
$\Omega_{\rm Bog}^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+c_{s}^{2}k^{2},$ (13)
where the sound speed $c_{s}^{2}=\frac{g_{\rm 1D,0}n_{z}}{m}$. The derivation of Eq. (11) from the linearized Eqs. (10a) and (10b) employs two key approximations to arrive at the standard form of the damped Mathieu equation. First, the weak damping approximation ($\gamma\ll\omega,\Omega_{k}$) is applied, whereby terms of order $\gamma^{2}$ and higher are neglected. This is consistent with the phenomenological inclusion of dissipation and justifies the retention of only the first-order term $2\gamma\dot{U}$. Second, we consider the long-wavelength, low-temperature regime where the thermal energy scale satisfies $\beta^{-1}\ll\epsilon_{k}$. This ensures that the correction from the logarithmic nonlinearity remains small compared to the quantum pressure term, allowing us to cleanly isolate the parametric driving term. Under these physically reasonable conditions, the dynamics of the perturbation amplitude $U$ is faithfully described by the Mathieu equation, which captures the interplay between parametric resonance and Jeans instability.
Figure 1 illustrates the collective excitation spectrum of SGBECs. At large $k$ ($k\xi\gg 1$, where $\xi=\hbar/\sqrt{2mg_{\rm 1D,0}n_{z}}$ is the healing length), $\Omega_{k}$ approaches the Bogoliubov dispersion. The critical wavenumber $k_{J}$, indicated by the vertical dashed line in Fig. 1, marks the onset of Jeans instability and is defined by the condition $\Omega_{k}^{2}=0$. For $k<k_{J}$ (specifically, $k_{J}\xi\equiv\sqrt{\sqrt{1+\Omega^{2}_{J}/\Omega^{2}_{0}}-1}$, where $\Omega_{0}=g_{\rm 1D,0}n_{z}/\hbar$), the Jeans instability causes $\Omega_{k}^{2}<0$, indicating exponential growth. This critical wavenumber $k_{J}$ defines the quantum Jeans scale Chavanis (2020).
Figure 1: (Color online) Collective excitation spectrum of SGBECs. Parameters: $\Omega_{0}=g_{\rm 1D,0}n_{z}/\hbar$, $\xi=\hbar/\sqrt{2mg_{\rm 1D,0}n_{z}}$. The vertical dashed line indicates the critical wavenumber $k_{J}$ where $\Omega_{k}^{2}=0$. The region ($k<k_{J}$) indicates Jeans instability with $\Omega_{k}^{2}<0$.
Eq. (11) is rendered dimensionless via $\tau=\omega t/2$:
$\frac{{\rm d}^{2}U}{{\rm d}\tau^{2}}+\bar{\gamma}\frac{{\rm d}U}{{\rm d}\tau}+\left[\lambda-2q\cos(2\tau)\right]U=0,$ (14)
with parameters
$\lambda=\frac{4\Omega_{k}^{2}}{\omega^{2}},\quad q=-\frac{2\alpha g_{\rm 1D,0}n_{z}}{m}\frac{k^{2}}{\omega^{2}},$ (15)
and $\bar{\gamma}\equiv 4\gamma/\omega$.
In the dissipationless limit ($\gamma=0$), Eq. (14) is reduced to the canonical Mathieu equation, whose stability properties are well-established McLachlan (1947); Abramowitz and Stegun (1965); Wang and Guo (1989). The parameter $\lambda$ encodes the underlying excitation spectrum of the self-gravitating quantum fluid. We analyze its phase diagram and derive the Faraday wavevector in the following section.